How do I find the function of f to find f(2), f(3), f(4)? Do not have a graph
Answer by aliateach(10)

to find a the output of a function you have two options
1)Using Algebraic formulas
2)graphs
Search for the formula
e.g
f(x) = x+1
then substitute for every x value in the expression
f(1) = 1 +1 = 2
f(2) = 2 +1 = 3
another example
f(x) = x^2 -1
f(1) = 1^1 -1
f(1) = 0
f(2)=2^2 - 1
f(2) = 4-1
f(2)=3